The demand curve for a product is given by the equation Qa 60 4 P And the supply curve for the product is given by the equation Qs = 3 P - 10. (P is measured in S) The absolute value of price-elasticity of demand at the market equilibrium price and quantity is (a) 0.15 (b) 0.75 (d) 2.0
Answer #1

Ans: d) 2.0

Explanation:

At equilibrium,

Quantity demanded = Quantity supplied

60 - 4P = 3P -10

7P = 70

P = 70/ 7 = 10

Quantity demand = 60 - 4P

= 60 - (4*10)

=60 - 40

=20

PED = (\Delta Q / \Delta P)* ( P / Q )

= - 4*(10/20)

= -4 * 0.5

=-2.0

The absolute value of PED is 2.0